OpenStudy (anonymous):

Consider the function f(x) = (√2x+1/4x-3) . Determine the largest possible domain of f(x) .Write your answer in interval notation.

OpenStudy (anonymous):

What is your function \[\frac{ \sqrt{2x+1} }{ 4x-3 }\] or \[\frac{\sqrt{2x}+1}{4x-3}\]

OpenStudy (anonymous):

√(2x+1/4x-3)

OpenStudy (anonymous):

all the stuff is under square root

OpenStudy (anonymous):

You means this function ?\[\sqrt{\frac{ 2x+1 }{ 4x-3 }}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Domain is\[(-\infty , -\frac{1}{2} ] \cup (\frac{3}{4},+\infty)\]